# Scalable Bayesian Optimization via Focalized Sparse Gaussian Processes

### v1

@@ 1 Introduction
Bayesian Optimization (BO) is a powerful approach for solving black-box optimization problems, demonstrating notable success in hyperparameter tuning (Snoek et al., 2012), reinforcement learning (Calandra et al., 2016; Chen et al., 2018), and scientific discovery (Gómez-Bombarelli et al., 2018). The efficacy of BO is attributed to its ability to model the unknown objective function using a surrogate model and to strategically select the next sample position by optimizing an acquisition function. Among the surrogate models, Gaussian Processes (GPs) (Rasmussen, 2003) are usually favored due to their flexibility and robust uncertainty quantification. However, the computation of the posterior GP covariance matrix scales as $\mathcal{O}(n^{3})$ with the number of data points $n$, which can severely restrict the applicability of BO in handling large datasets. This poses a significant challenge for real-world applications with high-dimensional and heterogeneous function landscapes such as those in robot control, which often necessitate a substantial amount of data to adequately explore the vast search space. To extend the scope of BO to accommodate larger datasets (from long-horizon online trials and/or pre-collected offline datasets (Trabucco et al., 2022)), it is imperative to employ surrogate models that offer enhanced computational efficiency.
Using sparse GP models is a popular method for reducing the computational cost of BO. Sparse GPs accomplish this by learning an approximation of the full GP, either by using a subset of data (Lawrence et al., 2002), ensemble of local models (Snelson and Ghahramani, 2005), or variational inference (Titsias, 2009). However, classical sparse GP models are typically tailored for regression tasks, and therefore are designed to fit to the entire function landscape. Given limited representational resources, the resulting posterior is likely to be overly smooth, which may negatively impact the performance of BO. This issue is exacerbated in the high-dimensional setting, in which accurately fitting the entire domain is a far more challenging task. As such, several works have proposed strategies to improve BO performance with sparse GP models by focusing promising regions (McIntire et al., 2016; Moss et al., 2023) or advanced sparse GP models (Jimenez and Katzfuss, 2023). However, most of their empirical evaluations are only conducted under large online sample setting in low-dimensional problems with fewer than 20 variables. It is unclear whether existing methods can be generalized to large offline data or high-dimensional setting.
@@ 2.1 Sparse Gaussian processes
Scaling Gaussian processes to large datasets is an important topic (Liu et al., 2020). It can be broadly divided into global approximation strategies and local approximation strategies.
Global sparse GPs perform distillation over the whole dataset to approximate the expensive full covariance matrix with a sparse representation. Several methods aim to choose a subset of representative training points from the whole dataset, and use the corresponding covariance matrix in place of the full covariance (Hayashi et al., 2020; Lawrence et al., 2002; Seeger, 2002; Keerthi and Chu, 2005). Sparse kernels aim at removing uncorrelated entries in the full covariance to obtain a compact matrix (Gneiting, 2002; Melkumyan and Ramos, 2009; Buhmann, 2001; Wendland, 2004). Sparse approximation methods use inducing variables to learn a low-rank representation of full covariance matrix (Quinonero-Candela and Rasmussen, 2005; Smola and Bartlett, 2000; Seeger et al., 2003; Csató and Opper, 2002; Snelson and Ghahramani, 2005; Titsias, 2009; Hensman et al., 2013; Csató and Opper, 2000; Wilson and Nickisch, 2015). Stochastic variational Gaussian process (SVGP) is a popular sparse GP method which employs variational inference to learn inducing variables and kernel hyperparameters jointly and enable training using stochastic gradient descent from mini-batch data (Hensman et al., 2013). Recently, nearest neighbor information has also been used to further improve the scalability of sparse GP over massive amount of data (Wu et al., 2022; Tran et al., 2021).
In contrast, local sparse GPs divide the entire dataset and employ local GPs trained from different data subsets to approximate the full GP. For a given test set, the prediction can be extracted from one of the local GPs (Kim et al., 2005; Datta et al., 2016), mixture of GPs (Yuksel et al., 2012; Masoudnia and Ebrahimpour, 2014) or product of GPs (Hinton, 2002; Cohen et al., 2020).
@@ 2.2 Scalable Bayesian optimization
Recent works have proposed modifications to sparse GPs for Bayesian optimization. Sparse GP has been used to determine the search region where local GPs are used to determine the next samples (Krityakierne and Ginsbourger, 2015). Weighted-update online Gaussian processes (WOGP) was developed to select a subset of training points to approximate high performing regions of the input space (McIntire et al., 2016). IMP-DPP is motivated by a similar observation and uses a weighted Determinantal Point Process to select training points as inducing variables for the SVGP (Moss et al., 2023). However, their proposed selection strategies require sequentially evaluating every training point, which can be computationally very expensive with large offline datasets. Combining SVGP with Thompson sampling has the same order of regret as standard Thompson sampling method (Vakili et al., 2021). Online variational conditioning (OVC) was proposed to efficiently conditioning SVGPs in an online setting, enabling using look-ahead acquisition functions (Maddox et al., 2021). Vecchia approximation of GP was also applied (Katzfuss et al., 2020) for Bayesian optimization, with improved performance compared to prior works (Jimenez and Katzfuss, 2023). A concurrent work (Maus et al., 2024) aims at improving the acquisition optimization performance based on target-aware Bayesian inference (Rainforth et al., 2020).
Besides sparse GPs, Neural network (Snoek et al., 2015; Shangguan et al., 2021) and random forest (Lakshminarayanan et al., 2016) can also be used as BO surrogate model to circumvent the cubic complexity of GP. Ensemble Bayesian optimization utilizes the addictive function structure and uses ensembles of addictive GPs in parallel to achieve scalability (Wang et al., 2018). Trust Region Bayesian optimization (TuRBO) and its variants uses exact GP to optimize over local regions, and employs a restart mechanism to achieve large number of evaluation, which is a representative line of works in high-dimensional Bayesian optimization (Eriksson et al., 2019; Wang et al., 2020; Eriksson and Poloczek, 2021). TuRBO can also be combined with sparse GP models to further enhance the scalability (Maus et al., 2022; Tautvaišas and Žilinskas, 2024).
@@ 3.1 Bayesian optimization
where $K$ is the covariance matrix between subscript inputs. With the posterior distribution given $D$, the next sample point is the maximum position of the acquisition function: $\boldsymbol{x}_{t+1}=\max_{\boldsymbol{x}\in\mathcal{X}}a(\boldsymbol{x}|\mathcal{M}_{t})$, where $\mathcal{M}_{t}$ is the GP model fitted on dataset collected at time step $t$. Common-used choice of $a$ includes upper confidence bound(UCB, (Srinivas et al., 2009)), expected improvement (EI, (Jones et al., 1998)) and Thompson sampling (TS, (Kandasamy et al., 2018)). The inner optimization problem is usually solved by grid search, evolutionary algorithms (Hansen, 2006), or gradient-based methods (Balandat et al., 2020). When the online sample budget is large, batch optimization is commonly used to evaluate multiple inputs in parallel (González et al., 2016).
@@ 3.2 Variational Gaussian process
Variational GP (Titsias, 2009) considers the joint latent prior
$\begin{split}p(\boldsymbol{f},\boldsymbol{u})=(\begin{bmatrix}\boldsymbol{f}\\ \boldsymbol{u}\end{bmatrix}\mid 0,\begin{bmatrix}K_{\boldsymbol{X}\boldsymbol{X}}&K_{\boldsymbol{X}\boldsymbol{Z}}\\ K_{\boldsymbol{Z}\boldsymbol{X}}&K_{\boldsymbol{Z}\boldsymbol{Z}}\end{bmatrix}),\end{split}$ (2)
@@
[…]
where $\boldsymbol{A}=K_{\boldsymbol{ZZ}}^{-1}K_{\boldsymbol{Z}\boldsymbol{X}}$. The variational parameters $\boldsymbol{Z},\boldsymbol{m},\boldsymbol{S}$ are optimized by maximizing the Evidence Lower Bound (ELBO) which can be written in the following formulation (Hensman et al., 2013):
@@
[…]
@@ 4 Focalized Gaussian Process for Bayesian Optimization
Prior studies about variational sparse GPs are mainly designed for regression tasks, where the goal is to fit global training data distribution. In Bayesian optimization, the next sample is determined by the predictive function distribution over test positions. Gradient-based and evolutionary-based acquisition function optimization methods employ local search from random starting points to find a local optimal of the acquisition function. Recent works also scale grid search-based optimization to high dimensional space by restricting the search space within local sub-regions (Eriksson et al., 2019; Wang et al., 2020). All the above procedure would benefit from an accurate estimation over sub-region of the input space. Therefore, a sensible way to improve BO performance is to allocate limited computational resources to obtain better prediction over specific search regions instead of the entire input domain.
@@ 4.1 Focalized evidence lower bound
where $\bar{k}_{ij}$ is the $(i,j)$-th entry of $[K_{\boldsymbol{X}\boldsymbol{X}}+\sigma I]^{-1}$. From eq. 6 we can observe that the mean estimation at $\boldsymbol{x}_{*}$ is a linear combination of observation $\boldsymbol{y}$ multiplied by $k(\boldsymbol{\boldsymbol{x}_{*}},\boldsymbol{x})$, and the reduction of variance is a quadratic form of the covariance between $\boldsymbol{x}_{*}$ and training points. Both estimation can be written as linear summations of constant values with kernel function as weight. As mentioned in prior works (Gramacy and Apley, 2015), data points far from the test positions have a vanishingly small influence on the predictive distribution with commonly used kernel functions. Utilizing this observation, we propose to weight the data likelihood term using the kernel function to focus training over points that contribute to the prediction of the search region:
@@
Suppose that we know the optimal point lies in some small sub-region of $\boldsymbol{X}$ that contains $N^{\prime}<<N$ training points. Corollary 19 in (Burt et al., 2020) shows that given a squared exponential kernel and some assumptions on the inducing point selection, for a fixed number of inducing points the KL-divergence upper bound scales super-quadratically in the number of training points. Hence, fitting locally can yield much tighter approximations than fitting globally (e.g. SVGP).
Next, we consider the impact of the KL approximation error on the optimization regret. Proposition 1 in (Burt et al., 2020) states that the gap between the means of the approximate and exact posteriors is upper bounded by $\mathcal{O}(\sigma\sqrt{\gamma})$, where $\gamma$ is an upper-bound on the approximation KL-divergence. This has an immediate impact on the regret - for example, when GP-UCB (Srinivas et al., 2009) is combined with sparse GPs, the confidence bounds must be enlargened by an additive $\sqrt{\gamma}$ factor to account for the approximation error. Because the regret bound scales with $\sqrt{\beta_{T}}$ where $\beta_{T}$ is the maximum confidence interval coefficient, having a large approximation error can arbitrarily scale the regret incurred by the algorithm. In order to achieve no additional regret order, the additional approximation error noise must be uniformly bounded (Assumption 4 in (Vakili et al., 2021)). Although focalized GP cannot guarantee a constant bound, it still directly reduces the regret of the algorithm, where we empirically investigate in Appendix B.1.
@@ 4.2 Bayesian optimization with focalized GP
At each BO iteration, FocalBO  iteratively optimizes the acquisition function over a progressively smaller search region via focalized acquisition function (FocalAcq) as shown in Algorithm 2. The first depth of acquisition optimization starts with the entire input space $\mathcal{X}$ with $\boldsymbol{l}=(1,\cdots,1)^{T}$ and $\boldsymbol{c}=(0.5,\cdots,0.5)^{T}$ (line 1). We train specific focalized GP base on the search region at each round of acquisition optimization (line 4-5). Our framework is compatible with any acquisition function that extracts instant posterior information from the GP and is optimized within pre-defined search region. After one round of acquisition function optimization, the search space length $\boldsymbol{l}$ is halved to focus on a smaller search region centered at current best position $\boldsymbol{x}_{\text{best}}$(line 6-7). In this way we can obtain a more accurate model for decision making, and also relieve the over-exploration problem when the problem dimension is high (Oh et al., 2018). One batch of inputs is proposed at each round of optimization, and the final decision is sampled from all proposed inputs via Softmax distribution over their corresponding acquisition function values (line 9). Our hierarchical optimization strategy enables collecting candidates from both global sparse estimation and local focalized prediction, achieving balance between exploration and exploitation with constrained computation power.
@@
1:  $\boldsymbol{l}\leftarrow(1,\cdots,1)^{T},\boldsymbol{c}\leftarrow(0.5,\cdots,0.5)^{T}$
@@
Our proposed framework is orthogonal to TuRBO-M (Eriksson et al., 2019), but bears some similarities in searching over multiple sub-regions and adaptively adjusting the search region. Our algorithm differs in that TuRBO-M constructs equal-sized trust regions and fits independent Exact GP using separated dataset, aiming at searching for different local optima in the search space. By contrast, the search region in FocalBO is constructed with different sizes to make decision based on both global and local information. Our framework allows data sharing across search regions, and the use of focalized GP helps to accurately estimate local region with limited representation. Additionally, FocalBO does not introduce extra hyperparameters. Finally, we demonstrate in Section 5 that TuRBO is complementary to FocalBO in optimizing high-dimensional problems.

### v2

@@ 1 Introduction
Bayesian Optimization (BO) is a powerful approach for solving black-box optimization problems, demonstrating notable success in hyperparameter tuning (snoek2012practical), reinforcement learning (calandra2016bayesian; chen2018bayesian), and scientific discovery (gomez2018automatic). The efficacy of BO is attributed to its ability to model the unknown objective function using a surrogate model and to strategically select the next sample position by optimizing an acquisition function. Among the surrogate models, Gaussian Processes (GPs) (rasmussen2003gaussian) are usually favored due to their flexibility and robust uncertainty quantification. However, the computation of the posterior GP covariance matrix scales as $\mathcal{O}(n^{3})$ with the number of data points $n$, which can severely restrict the applicability of BO in handling large datasets. This poses a significant challenge for real-world applications with high-dimensional and heterogeneous function landscapes such as those in robot control, which often necessitate a substantial amount of data to adequately explore the vast search space. To extend the scope of BO to accommodate larger datasets (from long-horizon online trials and/or pre-collected offline datasets (trabucco2022design)), it is imperative to employ surrogate models that offer enhanced computational efficiency.
Using sparse GP models is a popular method for reducing the computational cost of BO. Sparse GPs accomplish this by learning an approximation of the full GP, either by using a subset of data (lawrence2002fast), ensemble of local models (snelson2005sparse), or variational inference (titsias2009variational). However, classical sparse GP models are typically tailored for regression tasks, and therefore are designed to fit to the entire function landscape. Given limited representational resources, the resulting posterior is likely to be overly smooth, which may negatively impact the performance of BO. This issue is exacerbated in the high-dimensional setting, in which accurately fitting the entire domain is a far more challenging task. As such, several works have proposed strategies to improve BO performance with sparse GP models by focusing promising regions (mcintire2016sparse; moss2023inducing) or advanced sparse GP models (jimenez2023scalable). However, most of their empirical evaluations are only conducted under large online sample setting in low-dimensional problems with fewer than 20 variables. It is unclear whether existing methods can be generalized to large offline data or high-dimensional setting.
@@ 2.1 Sparse Gaussian processes
Scaling Gaussian processes to large datasets is an important topic (liu2020gaussian). It can be broadly divided into global approximation strategies and local approximation strategies.
Global sparse GPs perform distillation over the whole dataset to approximate the expensive full covariance matrix with a sparse representation. Several methods aim to choose a subset of representative training points from the whole dataset, and use the corresponding covariance matrix in place of the full covariance (hayashi2020random; lawrence2002fast; seeger2002pac; keerthi2005matching). Sparse kernels aim at removing uncorrelated entries in the full covariance to obtain a compact matrix (gneiting2002compactly; melkumyan2009sparse; buhmann2001new; wendland2004scattered). Sparse approximation methods use inducing variables to learn a low-rank representation of full covariance matrix (quinonero2005unifying; smola2000sparse; seeger2003fast; csato2002sparse; snelson2005sparse; titsias2009variational; hensman2013gaussian; csato2000sparse; wilson2015kernel). Stochastic variational Gaussian process (SVGP) is a popular sparse GP method which employs variational inference to learn inducing variables and kernel hyperparameters jointly and enable training using stochastic gradient descent from mini-batch data (hensman2013gaussian). Recently, nearest neighbor information has also been used to further improve the scalability of sparse GP over massive amount of data (wu2022variational; tran2021sparse).
In contrast, local sparse GPs divide the entire dataset and employ local GPs trained from different data subsets to approximate the full GP. For a given test set, the prediction can be extracted from one of the local GPs (kim2005analyzing; datta2016nearest), mixture of GPs (yuksel2012twenty; masoudnia2014mixture) or product of GPs (hinton2002training; cohen2020healing).
@@ 2.2 Scalable Bayesian optimization
Recent works have proposed modifications to sparse GPs for Bayesian optimization. Sparse GP has been used to determine the search region where local GPs are used to determine the next samples (krityakierne2015global). Weighted-update online Gaussian processes (WOGP) was developed to select a subset of training points to approximate high performing regions of the input space (mcintire2016sparse). IMP-DPP is motivated by a similar observation and uses a weighted Determinantal Point Process to select training points as inducing variables for the SVGP (moss2023inducing). However, their proposed selection strategies require sequentially evaluating every training point, which can be computationally very expensive with large offline datasets. Combining SVGP with Thompson sampling has the same order of regret as standard Thompson sampling method (vakili2021scalable). Online variational conditioning (OVC) was proposed to efficiently conditioning SVGPs in an online setting, enabling using look-ahead acquisition functions (maddox2021conditioning). Vecchia approximation of GP was also applied (katzfuss2020vecchia) for Bayesian optimization, with improved performance compared to prior works (jimenez2023scalable). A concurrent work (maus2024approximation) aims at improving the acquisition optimization performance based on target-aware Bayesian inference (rainforth2020target).
Besides sparse GPs, Neural network (snoek2015scalable; shangguan2021neural) and random forest (lakshminarayanan2016mondrian) can also be used as BO surrogate model to circumvent the cubic complexity of GP. Ensemble Bayesian optimization utilizes the addictive function structure and uses ensembles of addictive GPs in parallel to achieve scalability (wang2018batched). Trust Region Bayesian optimization (TuRBO) and its variants uses exact GP to optimize over local regions, and employs a restart mechanism to achieve large number of evaluation, which is a representative line of works in high-dimensional Bayesian optimization (eriksson2019scalable; wang2020learning; eriksson2021scalable). TuRBO can also be combined with sparse GP models to further enhance the scalability (maus2022local; tautvaivsas2024scalable).
@@ 3.1 Bayesian optimization
where $K$ is the covariance matrix between subscript inputs. With the posterior distribution given $D$, the next sample point is the maximum position of the acquisition function: $\boldsymbol{x}_{t+1}=\max_{\boldsymbol{x}\in\mathcal{X}}a(\boldsymbol{x}|\mathcal{M}_{t})$, where $\mathcal{M}_{t}$ is the GP model fitted on dataset collected at time step $t$. Common-used choice of $a$ includes upper confidence bound(UCB, (srinivas2009gaussian)), expected improvement (EI, (jones1998efficient)) and Thompson sampling (TS, (kandasamy2018parallelised)). The inner optimization problem is usually solved by grid search, evolutionary algorithms (hansen2006cma), or gradient-based methods (balandat2020botorch). When the online sample budget is large, batch optimization is commonly used to evaluate multiple inputs in parallel (gonzalez2016batch).
@@ 3.2 Variational Gaussian process
Variational GP (titsias2009variational) considers the joint latent prior
[…]
@@
$q(\boldsymbol{f})=\int p(\boldsymbol{f}\mid\boldsymbol{u})q(\boldsymbol{u})d\boldsymbol{u}=\mathcal{N}(\boldsymbol{f}\mid\boldsymbol{A}\boldsymbol{m},K_{\boldsymbol{X}\boldsymbol{X}}-\boldsymbol{A}^{T}(K_{\boldsymbol{Z}\boldsymbol{Z}}-\boldsymbol{S})\boldsymbol{A}),$ (3)
where $\boldsymbol{A}=K_{\boldsymbol{ZZ}}^{-1}K_{\boldsymbol{Z}\boldsymbol{X}}$. The variational parameters $\boldsymbol{Z},\boldsymbol{m},\boldsymbol{S}$ are optimized by maximizing the Evidence Lower Bound (ELBO) which can be written in the following formulation (hensman2013gaussian):
@@
where $\text{KL}[\cdot\parallel\cdot]$ is the KL divergence between two distributions. The ELBO breaks into a data likelihood term which factorized over training data, and a KL divergence term which can be computed in closed form. The factorization over data allows optimization via stochastic gradient descent (SGD), reducing the computational complexity to $\mathcal{O}(m^{3})$.
@@ 4 Focalized Gaussian Process for Bayesian Optimization
Prior studies about variational sparse GPs are mainly designed for regression tasks, where the goal is to fit global training data distribution. In Bayesian optimization, the next sample is determined by the predictive function distribution over test positions. Gradient-based and evolutionary-based acquisition function optimization methods employ local search from random starting points to find a local optimal of the acquisition function. Recent works also scale grid search-based optimization to high dimensional space by restricting the search space within local sub-regions (eriksson2019scalable; wang2020learning). All the above procedure would benefit from an accurate estimation over sub-region of the input space. Therefore, a sensible way to improve BO performance is to allocate limited computational resources to obtain better prediction over specific search regions instead of the entire input domain.
@@ 4.1 Focalized evidence lower bound
where $\bar{k}_{ij}$ is the $(i,j)$-th entry of $K_{\boldsymbol{X}\boldsymbol{X}}+\sigma I]^{-1}$. From eq. 6 we can observe that the mean estimation at $\boldsymbol{x}_{*}$ is a linear combination of observation $\boldsymbol{y}$ multiplied by $k(\boldsymbol{\boldsymbol{x}_{*}},\boldsymbol{x})$, and the reduction of variance is a quadratic form of the covariance between $\boldsymbol{x}_{*}$ and training points. Both estimation can be written as linear summations of constant values with kernel function as weight. As mentioned in prior works (gramacy2015local), data points far from the test positions have a vanishingly small influence on the predictive distribution with commonly used kernel functions. Utilizing this observation, we propose to weight the data likelihood term using the kernel function to focus training over points that contribute to the prediction of the search region:
@@
Suppose that we know the optimal point lies in some small sub-region of $\boldsymbol{X}$ that contains $N^{\prime}<<N$ training points. Corollary 19 in (burt2020convergence) shows that given a squared exponential kernel and some assumptions on the inducing point selection, for a fixed number of inducing points the KL-divergence upper bound scales super-quadratically in the number of training points. Hence, fitting locally can yield much tighter approximations than fitting globally (e.g. SVGP).
Next, we consider the impact of the KL approximation error on the optimization regret. Proposition 1 in (burt2020convergence) states that the gap between the means of the approximate and exact posteriors is upper bounded by $\mathcal{O}(\sigma\sqrt{\gamma})$, where $\gamma$ is an upper-bound on the approximation KL-divergence. This has an immediate impact on the regret - for example, when GP-UCB (srinivas2009gaussian) is combined with sparse GPs, the confidence bounds must be enlargened by an additive $\sqrt{\gamma}$ factor to account for the approximation error. Because the regret bound scales with $\sqrt{\beta_{T}}$ where $\beta_{T}$ is the maximum confidence interval coefficient, having a large approximation error can arbitrarily scale the regret incurred by the algorithm. In order to achieve no additional regret order, the additional approximation error noise must be uniformly bounded (Assumption 4 in (vakili2021scalable)). Although focalized GP cannot guarantee a constant bound, it still directly reduces the regret of the algorithm, where we empirically investigate in Appendix B.1.
@@ 4.2 Bayesian optimization with focalized GP
At each BO iteration, FocalBO  iteratively optimizes the acquisition function over a progressively smaller search region via focalized acquisition function (FocalAcq) as shown in Algorithm 2. The first depth of acquisition optimization starts with the entire input space $\mathcal{X}$ with $\boldsymbol{l}=(1,\cdots,1)^{T}$ and $\boldsymbol{c}=(0.5,\cdots,0.5)^{T}$ (line 1). We train specific focalized GP base on the search region at each round of acquisition optimization (line 4-5). Our framework is compatible with any acquisition function that extracts instant posterior information from the GP and is optimized within pre-defined search region. After one round of acquisition function optimization, the search space length $\boldsymbol{l}$ is halved to focus on a smaller search region centered at current best position $\boldsymbol{x}_{\text{best}}$(line 6-7). In this way we can obtain a more accurate model for decision making, and also relieve the over-exploration problem when the problem dimension is high (oh2018bock). One batch of inputs is proposed at each round of optimization, and the final decision is sampled from all proposed inputs via Softmax distribution over their corresponding acquisition function values (line 9). Our hierarchical optimization strategy enables collecting candidates from both global sparse estimation and local focalized prediction, achieving balance between exploration and exploitation with constrained computation power.
@@
[…]
@@
Our proposed framework is orthogonal to TuRBO-M (eriksson2019scalable), but bears some similarities in searching over multiple sub-regions and adaptively adjusting the search region. Our algorithm differs in that TuRBO-M constructs equal-sized trust regions and fits independent Exact GP using separated dataset, aiming at searching for different local optima in the search space. By contrast, the search region in FocalBO is constructed with different sizes to make decision based on both global and local information. Our framework allows data sharing across search regions, and the use of focalized GP helps to accurately estimate local region with limited representation. Additionally, FocalBO does not introduce extra hyperparameters. Finally, we demonstrate in Section 5 that TuRBO is complementary to FocalBO in optimizing high-dimensional problems.